A 1000 kg truck is parked on a bridge 100 m long at a distance of 25 m from one end of
the bridge. The mass of the bridge is 200 kg m−1
. Calculate the reaction forces at the
supports at the two ends of the bridge. (Take g = 10 ms−2
)
1
Expert's answer
2019-03-25T08:48:57-0400




"\sum M_A=0\\
R_BL=W_{\rm{truck}}d+W_{\rm{bridge}}L/2"

"R_B=\frac{W_{\rm{truck}}d+W_{\rm{bridge}}L/2}{L}=\frac{1000\times 10\times 25+200\times 10\times 50}{100}"

"R_B=3500\:\rm{N}"

"\sum M_B=0\\
R_AL=W_{\rm{truck}}(L-d)+W_{\rm{bridge}}L/2"

"R_A=\frac{W_{\rm{truck}}(L-d)+W_{\rm{bridge}}L/2}{L}=\frac{1000\times 10\times 75+200\times 10\times 50}{100}"

"R_A=8500\:\rm{N}"